# DGLA Actions: An Application in GR

### v1

@@ Acknowledgements
I thank my collaborators on related projects Filip Dul, Surya Raghavendran, and Changsun Choi. I am also appreciative of several discussions with Peter Hintz and Justin Corvino regarding the gluing of spacetimes. The author is supported by the Simons Foundation under Travel Support/Collaboration 9966728.
@@ 2. DGLAs and Actions
    A graded Lie algebra $(\mathfrak{g}^{\ast},[-,-])$ over $\mathbb{K}$, and
@@
    A $\mathbb{K}$-linear map of degree one $d\colon\mathfrak{g}^{\ast}\to\mathfrak{g}^{\ast+1}$.
@@
    The map $d$ is a derivation of the bracket, i.e., for all $X,Y\in\mathfrak{g}^{\ast}$
@@
A (strict) map of dglas $\varphi\colon\mathfrak{g}^{\ast}\to\mathfrak{h}^{\ast}$ is a degree zero linear map which commutes with differentials and preserves brackets. General references for differential graded Lie algebras, their maps, and their actions (and the extension to $L_{\infty}$ algebras) are the book of Manetti [12] and Appendix A of the second volume of the tome by Costello and Gwilliam [5].
@@
Let $\mathfrak{g}^{\ast}$ and $\mathfrak{h}^{\ast}$ be differential graded Lie algebras. An action of $\mathfrak{g}^{\ast}$ on $\mathfrak{h}^{\ast}$ is the structure of a differential graded Lie algebra on $\mathfrak{g}^{\ast}\oplus\mathfrak{h}^{\ast}$ such that there is a short exact sequence
$0\to\mathfrak{h}^{\ast}\xrightarrow{\;\;\psi\;\;}\mathfrak{g}^{\ast}\oplus\mathfrak{h}^{\ast}\xrightarrow{\;\;\phi\;\;}\mathfrak{g}^{\ast}\to 0$
@@
Given a dgla $(\mathfrak{g}^{\ast},d,[-,-])$, it acts on itself (considered as a dgla) via the adjoint action. Indeed, the dgla structure on $\mathfrak{g}^{\ast}\oplus\mathfrak{g}^{\ast}$ is given by
@@ 2.1. Actions on Classical Field Theories
Given a classical field theory $(\mathcal{E},S)$, its covariant phase space is pre-symplectic and an action by a Lie algebra should be by symplectic vector fields. If we are interested in realizing symmetries as observables, e.g., as in Noether’s Theorem, then we should actually ask that our action factors through the Lie subalgebra of Hamiltonian vector fields. [^2] The notion of Hamiltonian vector field indeed makes sense in the pre-symplectic setting. A vector field $X$ is Hamiltonian if there exists a function $f$ such that $i_{X}\omega=df$.
The BV formalism provides a resolution of the (functions on reduced the) phase space. In the BV formalism, there is a natural (shifted) symplectic structure and corresponding Poisson bracket. The condition for our symmetry action now becomes that $\mathfrak{g}$ acts via symplectic vector fields which commute with the vector field $\{S,-\}$. One mathematical formulation of actions/symmetries is provided in Part 3 of Volume 2 of Costello–Gwilliam [5]. Moreover, Costello and Gwilliam reformulate actions in terms of equivariant actions, see Definition 12.2.12 of [5]. This equivariant action is the BV incarnation of our Definition 2.2 from above. In our application below, we will simply describe the explicit equivariant action. Note that for completeness, we should verify the equivariant classical master equation in the BV formulation of our theory; we do not do this here, details are given in [3].
@@ 3. Application: Geometry on a Neighborhood of Infinity via an Equivariant Action
As $\bf e$} is an isomorphism, the first equation is equivalent to $d_{\omega}{\bf e}=0$ and there is a unique metric connection satisfying this condition (with respect to $g_{\bf e}$), the Levi–Civita connection. The second equation then becomes the $\Lambda$-vacuum Einstein equation for the metric $g_{\bf e}$, see Remark 3.2 of [1].
@@ 3.1.2. Spherical Symmetry
Appendix A Further Remarks and Details

### v2

@@ 1. Introduction
In Section 2, we define dglas and their actions. We also indicate how a dgla acts on a classical field theory. In Section 3, we use dgla actions to constrain spacetime geometry. That is, we have a dgla which acts on (classical) Palatini–Cartan Theory and determines the Killing algebra outside of a compact region.
@@ Acknowledgements
I thank my collaborators on related projects Filip Dul, Surya Raghavendran, and Changsun Choi. I am also appreciative of several discussions with Peter Hintz and Justin Corvino regarding the gluing of spacetimes. Finally, I thank the anonymous referee for several helpful suggestions. The author is supported by the Simons Foundation under Travel Support/Collaboration 9966728.
@@ 2. DGLAs and Actions
    A graded Lie algebra $(\mathfrak{g}^{\bullet},[-,-])$ over $\mathbb{K}$, and
@@
    A $\mathbb{K}$-linear map of degree one $d\colon\mathfrak{g}^{\bullet}\to\mathfrak{g}^{\bullet+1}$.
@@
    The map $d$ is a derivation of the bracket, i.e., for all $X,Y\in\mathfrak{g}^{\bullet}$
@@
A (strict) map of dglas $\varphi\colon\mathfrak{g}^{\bullet}\to\mathfrak{h}^{\bullet}$ is a degree zero linear map which commutes with differentials and preserves brackets. General references for differential graded Lie algebras, their maps, and their actions (and the extension to $L_{\infty}$ algebras) are the book of Manetti [14] and Appendix A of the second volume of the tome by Costello and Gwilliam [6].
@@
Let $\mathfrak{g}^{\bullet}$ and $\mathfrak{h}^{\bullet}$ be differential graded Lie algebras. An action of $\mathfrak{g}^{\bullet}$ on $\mathfrak{h}^{\bullet}$ is the structure of a differential graded Lie algebra on $\mathfrak{g}^{\bullet}\oplus\mathfrak{h}^{\bullet}$ such that there is a short exact sequence
$0\to\mathfrak{h}^{\bullet}\xrightarrow{\;\;\psi\;\;}\mathfrak{g}^{\bullet}\oplus\mathfrak{h}^{\bullet}\xrightarrow{\;\;\phi\;\;}\mathfrak{g}^{\bullet}\to 0$
@@
Given a dgla $(\mathfrak{g}^{\bullet},d,[-,-])$, it acts on itself (considered as a dgla) via the adjoint action. Indeed, the dgla structure on $\mathfrak{g}^{\bullet}\oplus\mathfrak{g}^{\bullet}$ is given by
@@ 2.1. Actions on Classical Field Theories
Given a classical field theory $(\mathcal{E},S)$, a symmetry action by a Lie algebra should be by symplectic vector fields on the space of classical solutions. If we are interested in realizing symmetries as observables, e.g., as in Noether’s Theorem, then we should actually ask that our action factors through the Lie subalgebra of Hamiltonian vector fields.
The BV formalism provides a resolution of the (functions on) the Euler–Lagrange locus. This resolution is with respect to natural gauge symmetries, we may wish to consider additional symmetries and their associated conserved currents/charges. In the BV formalism, there is a natural (shifted) symplectic structure and corresponding Poisson bracket. The condition for our symmetry action now becomes that $\mathfrak{g}$ acts via symplectic vector fields which commute with the vector field $\{S,-\}$.
One mathematical formulation of actions/symmetries is provided in Part 3 of Volume 2 of Costello–Gwilliam [6]. Moreover, Costello and Gwilliam reformulate actions in terms of equivariant actions, see Definition 12.2.12 of [6]. This equivariant action is the BV incarnation of our Definition 2.2 from above. The rough idea underlying Costello–Gwilliam’s approach is to write every theory as a family of (deformed) BF or Chern–Simons theories. As such, any classical field theory is itself presented by a Lie theoretic object. The further action of $\mathfrak{g}$ on our theory can be rephrased in terms of a certain Maurer–Cartan element in an auxiliary “action dgla”; this method is standard and simply an extension of writing a map between Lie algebras in terms of their associated Chevalley–Eilenberg cochains. This Mauer–Cartan element then determines an action functional on an extended space of fields: the equivariant action functional. This process is reversible, so the symmetry action and the equivariant action functional are equivalent data.
In our application below, we will simply describe the explicit equivariant action. Note that for completeness, we should verify the equivariant classical master equation in the BV formulation of our theory; we do not do this here, details are given in [4].
@@ 3. Application: Geometry on a Neighborhood of Infinity via an Equivariant Action
As is explained in Section 6.1 of [17], the first equation shows that the covariant derivative $d_{\omega}$ (and hence $\omega$ itself) is determined by the tetrad. There is a unique torsion free, metric connection with respect to the pullback metric $g_{\bf e}$, its Levi–Civita connection. It then follows that the second equation can be written in terms of the Riemann curvature tensor and upon contraction with the tetrad recovers Einstein’s field equations.
@@ 3.1.2. Spherical Symmetry
Appendix A Fundamental Concepts/Results